Rewrite each equation in intercept form. Show your steps.
Equation: y= 5+3(x-4)
- How do I do this, and what's the answer?

Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

Rewrite each equation in intercept form. Show your steps.
Equation: y= 5+3(x-4)
x-intercept = ?
Let y = 0 and solve for "x":
5 + 3x-12 = 0
3x = 7
x = 7/3
---------------------------
y-intercept = ?
Let x = 0 and solve for "y":
y = 5+ 3(0-4)
y = 5-12
y = -7
--------
Intercept Form:
x/(7/3) + y/(-7) = 1
---
3x/7 - y/7 = 1
===================
Cheers,
Stan H.

Answer by rfer(16322)

y=5+3x-12
y=3x-7